Calculate the mass of gold as -3000J of energy is released as melted gold cools from 1064 ⁰C to room temperature (23⁰C) _ grams ( 2 Decimals) (Csp Au= 0.129)
1
Expert's answer
2019-05-14T03:40:47-0400

Solution.

"Q = c \times m \times (t2-t1)"

"m = \frac{Q}{c\times(t2-t1)}""m = \frac{-3000}{0.129\times10^3\times(23-1064)}"

m = 0.02234 kg = 22.34 g

Answer:

m = 22.34 g